Write the quadratic equation whose roots are 3 and 2, and whose leading coefficient is 3.
Answer by rothauserc(4718)

(3x-6) * (x-3) = 0
:
3x^2 -15x +18 = 0
:
Note that 3x-6 = 0 and x = 2, x-3 = 0 and and x=3
: